in a H-atom, the e- is making 6.6*10615 rev/s around the nucleus in an orbit of radius 0.528 Angstrom. the magnetic moment(Am^2) will be 1*10-10. plz explain................

Grade:12th pass

1 Answers

jyoti dash
8 years ago
M=IA where M is magnetic field, I is current and A is area
so as we know that A=\pir2 where r is the radius
I=e/t and t=1/f where e is charge t=time and f=frequency
substitute the quantity to get the result